OpenStudy (anonymous):

solve each system by substitution. -x+3y=-10 -3x+y=-22

OpenStudy (anonymous):

Basically what you are going to do is to simplify one equation to figure out what x is. For example... \[-x + 3y = -10\]\[-x = -10 - 3y\]\[x = 10 + 3y\]Now we know what x is equivalent to. So we just substitute x in the other equation... \[-3(10 + 3y) + y = -22\]\[-30 - 9y + y = -22\]\[-30 - 8y = -22\]\[-8y = 8\]\[y = -1\]Now that we know what y equals we can substitute that pack into the first equation to find the value of x... \[x = 10 + 3(-1)\]\[x = 10 - 3\]\[x = 7\] So now we know that our answer is.... \[x = 7\]\[y = -1\]